From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu May 15 2003 - 00:14:59 EDT
------------------------------------------------------------ Date: Thu May 15 00:14:59 2003, Shift: 16:00-24:00 Supervisor: R. Debbe, with: Log book#14 pages 110 - 110 Shift Summary: ------------------------------------------------------------ The shift started with a good store. then things got bad, several stores were lost because of radiation monitor interlocks, test to see if STAR and PHENIX get more collitions if BRAHMS and PHOBOS are off. Anyway, around 22:30 a good store was achieved and we are taking data happily. We collected ~10K trigger 2 in 1/1A at 4 degrees ------------------------------------------------------------
This archive was generated by hypermail 2.1.5 : Thu May 15 2003 - 00:19:39 EDT